# Erased, But Not Forgotten: Erased Rectified Flow Transformers Still Remain Unsafe Under Concept Attack

### v2

@@ Appendix I Looking Forward: Toward Robust Benchmarks and Safer Erasure Strategies in Rectified Flow Models
Our study positions concept attack not as an end in itself, but as a diagnostic instrument for understanding the limits of concept erasure in Flux. By showing that erased concepts can still be reliably reactivated under multiple state-of-the-art defenses, we expose fundamental weaknesses in current approaches. This finding underscores that erasure today is less a permanent solution than a fragile suppression of localized attention signals. Figure 13 provides a visualization of attack success rates on the “nudity” concept across different erasure methods, demonstrating that most methods only achieve surface-level suppression and leave deep conceptual residues prone to reactivation, with CP (Chavhan et al., 2025) emerging as the strongest yet still imperfect defense.
@@
LLM Usage Statement
Large Language Models (LLMs) are used in this work only for grammar checking and light language refinement of certain parts of the Introduction and Conclusion sections to improve readability and presentation quality. All core contributions, including the design of the method, theoretical formulations, experimental setup, and analysis are independently conceived and implemented by the authors without the involvement of LLMs. The full implementation code, experimental scripts, and datasets are prepared entirely by the authors, ensuring the accuracy and reproducibility of the reported results. The authors take full responsibility for the limited use of LLMs in language polishing and for all claims made in this paper.

### v3

@@ Appendix I Looking Forward: Toward Robust Benchmarks and Safer Erasure Strategies in Rectified Flow Models
Our study positions concept attack not as an end in itself, but as a diagnostic instrument for understanding the limits of concept erasure in Flux. By showing that erased concepts can still be reliably reactivated under multiple state-of-the-art defenses, we expose fundamental weaknesses in current approaches. This finding underscores that erasure today is less a permanent solution than a fragile suppression of localized attention signals. Figure 14 provides a visualization of attack success rates on the “nudity” concept across different erasure methods, demonstrating that most methods only achieve surface-level suppression and leave deep conceptual residues prone to reactivation, with CP [6] emerging as the strongest yet still imperfect defense.